length of two sides of a right triangle are giving calculate the unknown lenght
1.a=15ft., b=20ft.,c=?
2.a=7mi.,b=?.,c=25mi

Answer by checkley77(12569)

1.a=15ft., b=20ft.,c=?
15^2+20^2=C^2
225+400=C^2
625=C^2
C=SQRT625
C=25 ANS.
2.a=7mi.,b=?.,c=25mi
7^2+B^2=25^2
49=B^2=625
B^2=625-49
B^2=576
B=SQRT576
B=24 ANS.